OpenStudy (ambi):

Which one of the following demonstrates the greatest acceleration? A. car going from 0 m/s to 100 m/s in 50 seconds B. small marble rolling from 10m/s to 20 m/s in 1 second C. 2,000-lb elephant walking at a steady speed of 5 m/s D. roller coaster moving uphill from 1 m/s to 2 m/s in 1 second

OpenStudy (ambi):

@Rushwr :)

OpenStudy (rushwr):

What do u think ?

OpenStudy (ambi):

C?

OpenStudy (rushwr):

No steady state means no acceleration , u go in the same speed

OpenStudy (rushwr):

Nop Acceleration = final velocity - initial velocity divided by the time taken

OpenStudy (rushwr):

the acceleration in A = 100-0 / 50 = 100/50 = 2ms^-1

OpenStudy (ambi):

It has got to be D then, right?

OpenStudy (rushwr):

Nop the answer is B Velocity in B = 20 -10/1 = 10/1 = 10ms^-1 Velocity in D = 2-1/1 = 1/1 = 1ms^-1

OpenStudy (rushwr):

Got it right?

OpenStudy (ambi):

thanks alot

OpenStudy (ambi):

yea

OpenStudy (rushwr):

oh no problem :)